OpenStudy (anonymous):

px + qy = r 2px - qy = 2r When solving this system of equations, x =

OpenStudy (anonymous):

Input in standard form the equation of the given line. The line with m = 1/2 and b = 3.

OpenStudy (noorik.):

try to sum the two equations (px+qy) + (2px-qy) = r + 2r then 3px = 3r px = r

OpenStudy (anonymous):

That can't be right because my options are p/r r/p 3r/2p

OpenStudy (noorik.):

no it's correct, because px=r then x = r/p

OpenStudy (anonymous):

Okay! Could you help with this one? Input in standard form the equation of the given line. The line that passes through (-1, -3) and (2, 1)

OpenStudy (noorik.):

try to find the slope between the two points, and it's a linear equation of the first degree

OpenStudy (anonymous):

I have no clue what any of that means.

OpenStudy (noorik.):

ok the slope law is \[slope (m) = \frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

Okay?

OpenStudy (noorik.):

\[m = \frac{ 1-(-3) }{ 2-(-1) } = \frac{ 4 }{ 3 }\]

OpenStudy (anonymous):

Last one.. can you help me with this? Input in standard form the equation of the given line. The line with m = 1/2 and b = 3.

OpenStudy (noorik.):

then to find the equation of the line \[\frac{ 4 }{ 3 } = \frac{ y-1 }{ x-2 }\] then 4x-8 = 3y-3 3y = 4x -5

OpenStudy (noorik.):

the last question y = x/2 +3